OpenStudy (anonymous):

how to write y=-x+3 in standard form

OpenStudy (texaschic101):

standard form : Ax + By = C y = -x + 3 -- add x to both sides x + y = -x + x + 3 -- combine like terms x + y = 3 ===> standard form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

i have some problem solving questions can u help

OpenStudy (anonymous):